what is the vertex form of y= - 1/3 + 2x + 7/3
Found 2 solutions by solver91311, Onmyoji.S:
Answer by solver91311(24713)



That is a straight line. There is no "vertex form"

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by Onmyoji.S(32)

y= - 1/3 + 2x + 7/3
y=6/3+2x
y=2+2x
then I don't know how to do it.
x=-b/2a